The Department of Labor and Employment (DOLE) found that 83% of the Filipinos think that having a college education is important to succeed in life. If a ramdom sample of seven(7) Filipinos is selected, find the probability that the exactly four (4) people will agree with that statement (use Binomial Distribution
1
Expert's answer
2020-06-22T17:25:15-0400

Let, X = the random variable denoting the number of people will agree with the given statement, X = 0, 1, ... , 7.


We have 83% of the Filipinos think that having a college education is important to succeed in life i.e. the probability of people agreeing with this statement is 83% = 0.83.


Then X ~ bin(n = 7, p = 0.83)


The p.m.f. of X is given by,


P(X = x) "= \begin{cases}
 \dbinom{7}{x}(0.83)^x(1-0.83)^{7-x} &\text{for } x=0,1,...,7 \\
 0 &\text{otherwise }
\end{cases}"


Now, the probability that the exactly 4 people will agree with that statement


= P(X = 4)


= "\dbinom{7}{4}(0.83)^4(1-0.83)^{7-4}"


= 0.0816 (rounded to 4 decimal places)


Answer: The probability that exactly four people will agree with the given statement is 0.0816.